OpenStudy (anonymous):

Could someone please help! find the product of (3-4 square root 7)^2 I will draw it to make it clear.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[(3-4\sqrt{7})^2=(3-4\sqrt{7})(3-4\sqrt{7})\]

OpenStudy (phi):

use FOIL just like it was (a+b)(x+y) see http://www.khanacademy.org/math/algebra/polynomials/v/multiplying-binomials